OpenStudy (el_arrow):

find the exact length of the parametric curve x= e^t cost y= e^t sint 0

OpenStudy (el_arrow):

\[x = e^t cost\]

OpenStudy (el_arrow):

\[y= e^t sint\]

OpenStudy (el_arrow):

hello

OpenStudy (freckles):

kinda confused you mention 4 variables...

OpenStudy (el_arrow):

but thats whats on my homework

OpenStudy (freckles):

or is t meant to be theta

OpenStudy (el_arrow):

i think that t means theta

OpenStudy (freckles):

that would be great then :p

OpenStudy (freckles):

so have you found dx/dt and dy/dt

OpenStudy (el_arrow):

do we get rid of the t's in this problem?

OpenStudy (freckles):

we don't need to

OpenStudy (el_arrow):

why not

OpenStudy (freckles):

\[L=\int\limits_0^{\frac{\pi}{5}}\sqrt{(\frac{dx}{dt})^2+(\frac{dy}{dt})^2} dt\]

OpenStudy (el_arrow):

the derivative of e^t is just e^t right?

OpenStudy (freckles):

yah!

OpenStudy (el_arrow):

okay so dx is e

OpenStudy (freckles):

\[x=e^t \cos(t) \\ \\ \text{ To find } \frac{dx}{dt} \text{ use product rule }\]

OpenStudy (el_arrow):

dx = -te^t sin(t)

OpenStudy (freckles):

\[\frac{dx}{dt}=\frac{d(e^t)}{dt} \cdot \cos(t)+\frac{d (\cos(t))}{dt} \cdot e^t\]

OpenStudy (freckles):

remember product rule is (uv)'=u times v'+v times u'

OpenStudy (el_arrow):

so dx= te^t cos(t) + - e^t sin(t)

OpenStudy (freckles):

i thought you said derivative of e^t was e^t where does the extra t come from

OpenStudy (el_arrow):

i typed too fast lol sorry

OpenStudy (freckles):

\[\frac{dx}{dt}=e^t \cos(t)-\sin(t)e^t \\ \frac{dx}{dt}=e^t(\cos(t)-\sin(t))\]

OpenStudy (freckles):

now we need to find dy/dt

OpenStudy (el_arrow):

dy = e^t sin(t) + e^t cos(t)

OpenStudy (freckles):

\[x=e^t \cos(t) \\ \frac{dx}{dt}=e^t(\cos(t)-\sin(t)) \\ y=e^tsin(t) \\ \frac{dy}{dt}=e^t \sin(t)+ e^t \cos(t) =e^t(\sin(t)+\cos(t))\] great now we got square these ugly thingys

OpenStudy (freckles):

\[(\frac{dx}{dt})^2=e^{2t}(\cos(t)-\sin(t))^2=e^{2t}[\cos^2(t)-2\cos(t)\sin(t)+\sin^2(t)] \\ (\frac{dx}{dt})^2=e^{2t}[1-\sin(2t)]\] you try to simplify dy/dt now

OpenStudy (el_arrow):

okay

OpenStudy (freckles):

by the way you might find using (a+b)^2=a^2+2ab+b^2 useful :)

OpenStudy (el_arrow):

okay

OpenStudy (el_arrow):

for the trig identities right?

OpenStudy (freckles):

I don't know what you mean

OpenStudy (el_arrow):

like for what you do for the (cos(t) - sin(t))^2

OpenStudy (freckles):

I used (a-b)^2=a^2-2ab+b^2

OpenStudy (el_arrow):

yep do the same to dy

OpenStudy (el_arrow):

i got e^2t[1+cos(2t)] for dy

OpenStudy (freckles):

let me check

OpenStudy (freckles):

\[(\frac{dy}{dt})^2=e^{2t}(\sin^2(t)+2\sin(t)\cos(t)+\cos^2(t)) \\ (\frac{dy}{dt})^2=e^{2t}(1+\sin(2t))\] the only thing I disagree about is you making 2sin(t)cos(t) equal to something other than sin(2t)

OpenStudy (freckles):

\[(\frac{dx}{dt})^2+(\frac{dy}{dt})^2=e^{2t}(1+\sin(2t))+e^{2t}(1-\sin(2t))\] combine like terms

OpenStudy (freckles):

distribute first then combine like terms

OpenStudy (freckles):

you will see somethign really cool happen :)

OpenStudy (el_arrow):

for the dy why is the 2sin(t)cos(t) equal to sin(2t)

OpenStudy (freckles):

it is always equal to that

OpenStudy (freckles):

it is a trig identity

OpenStudy (freckles):

sin(2t)=2sin(t)cos(t) cos(2t)=cos^2(t)-sin^2(t)

OpenStudy (el_arrow):

oh i need to remember that one

OpenStudy (freckles):

@El_Arrow are you still simplifying the thingy I asked you too?

OpenStudy (freckles):

\[(\frac{dx}{dt})^2+(\frac{dy}{dt})^2=e^{2t}(1+\sin(2t))+e^{2t}(1-\sin(2t)) \] this thingy combing like terms?

OpenStudy (el_arrow):

yes trying to

OpenStudy (freckles):

\[e^{2t} +e^{2t} \sin(2t)+e^{2t}-e^{2t}\sin(2t)\] === you know u+u=? and v-v=?

OpenStudy (el_arrow):

it would be 2e^2t

OpenStudy (freckles):

\[L=\int\limits\limits_0^{\frac{\pi}{5}}\sqrt{(\frac{dx}{dt})^2+(\frac{dy}{dt})^2} dt \] \[L=\int\limits\limits_0^{\frac{\pi}{5}}\sqrt{2e^{2t}} dt \]

OpenStudy (el_arrow):

well its